OpenStudy (anonymous):

Brenda’s scores on the first three of four 100-point science tests were 95, 92, and 89. What score does she need on her fourth science test to ensure an average score of at least 93? A) X ≥ 96 B) X ≤ 90 C) X ≥ 88 D) X ≤ 92

OpenStudy (whpalmer4):

her average of 4 evenly weighted tests needs to be \(\ge 93\) average of 4 tests = \(\dfrac{test1 + test2 + test3 + test4}{4}\) We know that tests1-3 were 95,92, 89, so plug in the numbers and see what you need... \[\frac{95+92+89+x}{4} \ge 93\] She better study :-)

OpenStudy (anonymous):

She'll make it with 100, not much less.